OpenStudy (jenniferjuice):

6 + x + y = x + 6 + y is an example of which algebraic property?

OpenStudy (jenniferjuice):

Distributive Property Associative Property of Addition Commutative Property of Addition Symmetric Property

OpenStudy (anonymous):

distributive property of multiplication

OpenStudy (anonymous):

just so no one is confused, my statement is not the answer

OpenStudy (zzr0ck3r):

commutative a+b = b+a associative (a+b) +c= a+(b+c) distributive c(a+b) = ca+cb symetrix a=b then b = a

OpenStudy (zzr0ck3r):

Oh I c

OpenStudy (zzr0ck3r):

I don't think there is a distributive property of addition on the real line so that is redundant @completeidiot

OpenStudy (anonymous):

true but if they're asking a question like this, they must not know what the distribution property is and how it doesnt apply to addition statements

OpenStudy (zzr0ck3r):

its actually the distribution over addition

OpenStudy (zzr0ck3r):

it only applies to addition statements a(bc) does not equal abac

OpenStudy (anonymous):

pure addition statements better?

OpenStudy (zzr0ck3r):

so multiplication distributes over addition

OpenStudy (zzr0ck3r):

on the real line...

OpenStudy (zzr0ck3r):

@jenniferjuice given my first comment can you get the answer?

OpenStudy (jenniferjuice):

symmetix ...?

OpenStudy (jenniferjuice):

@zzr0ck3r

OpenStudy (anonymous):

no

OpenStudy (anonymous):

commutative property of addition states that you can add them in any order associative property of addition states that you can group them in any way, or in other words add the first 2 numbers first then add the third or add the last 2 numbers first then add the first

OpenStudy (anonymous):

associative property will always have the parenthesis